# ROCKET-1: Master Open-World Interaction with Visual-Temporal Context Prompting

### v1

@@ 4.4 What Matters for ROCKET-1?
We conduct ablation studies on short-horizon tasks of Minecraft Interaction benchmark: “Hunt right sheep (
@@
Interaction-Type Information Fusion Location
| Interaction Type Fusion |
Hunt ( |
Mine ( |
| in transformer layer | ||
| in visual backbone |
@@ 5 Related Works
The task instruction of a multi-task policy directly determines the types of tasks the policy can solve. Most current approaches (Brohan et al., 2022, 2023; Lynch et al., 2023; Cai et al., 2023a; Huang et al., 2023) use natural language to describe task details and collect large amounts of text-demonstration data pairs to train a language-conditioned policy for interaction with the environment. Although natural language can express a wide range of tasks, it struggles to represent spatial relationships effectively. Additionally, gathering text-annotated demonstration data is costly, limiting the scalability of these methods. Alternatives, such as Lifshitz et al. (2023); Majumdar et al. (2022); Sundaresan et al. (2024), use images to drive goal-conditioned policies, typically learning through hindsight relabeling in a self-supervised manner. While this reduces the need for annotated data, future images are often insufficiently expressive, making it difficult to capture detailed task execution processes. Methods like Cai et al. (2023b); Jang et al. (2022) propose using reference videos to describe tasks, offering strong expressiveness but suffering from ambiguity, which may lead to inconsistencies between policy interpretation and human understanding, raising safety concerns. Gu et al. (2023) suggests representing tasks with rough robot arm trajectories, enabling novel task completion but only in fully observable environments, limiting its applicability in open-world settings. CLIPort (Shridhar et al., 2022), which addresses pick-and-place tasks by controlling the robot’s start and end positions using heatmaps, bears some resemblance to our proposed visual-temporal context prompting method. However, CLIPort focuses solely on the pick-and-place task solutions in a fully observable environment.
@@
Minecraft offers a highly open sandbox environment with complex tasks and free exploration, ideal for testing AGI’s adaptability and long-term planning abilities. Its rich interactions and dynamic environment simulate real-world challenges, making it an excellent testbed for AGI. One line of research focuses on low-level control policies in Minecraft. Baker et al. (2022) annotated a large YouTube Minecraft video dataset with actions and trained the first foundation agent in the domain using behavior cloning, but it lacks instruction-following capabilities. Cai et al. (2023a) employs a goal-sensitive backbone and horizon prediction module to enhance multi-task execution in partially observable environments, but it only solves tasks seen during training. Fan et al. (2022) fine-tunes a vision-language alignment model MineCLIP using YouTube video data, and incorporates it into a reward shaping mechanism for training a multi-task agent, though task transfer still requires extensive environment interaction. Lifshitz et al. (2023) uses hindsight-relabeling to learn an image-goal-conditioned policy and aligns image and text spaces via MineCLIP, but this approach is limited to short-horizon tasks. Cai et al. (2023b) uses reference videos as task representations and trains a video-conditioned policy through a latent variable model, but generating environment-consistent reference videos remains a challenge. Another research focus integrates vision-language models for long-horizon task planning in Minecraft (Yuan et al., 2023; Wang et al., 2024b; Qin et al., 2023; Zheng et al., 2023). DEPS (Wang et al., 2023b), the first to apply large language models in Minecraft, uses a four-step process to decompose tasks, achieving the diamond mining challenge with minimal training. Voyager (Wang et al., 2023a) highlights LLM-based agents’ autonomous exploration and skill-learning abilities. Jarvis-1 (Wang et al., 2023c) extends DEPS with multimodal memory, improving long-horizon task success rates by recalling past experiences. OmniJarvis (Wang et al., 2024a) learns a behavior codebook using self-supervised methods to jointly model language, images, and actions. MineDreamer (Zhou et al., 2024) fine-tunes VLMs and a diffusion model to generate goal images for task execution, though it faces challenges with image quality and consistency.
6 Conclusions
@@
In this paper, we propose a novel hierarchical agent architecture to address the challenges of open-world interaction. To bridge spatial information gaps, we introduce a visual-temporal context prompting method that communicates intent between the high-level reasoner and the low-level policy. We develop ROCKET-1, an object-segmentation-conditioned policy responsible for real-time interaction with objects in the environment. This architecture can also leverage SAM-2 plug-and-play to enhance the policy’s object-tracking capabilities. Extensive experiments in Minecraft demonstrate that our approach successfully harnesses the powerful visual-language reasoning abilities of VLMs for embodied decision-making, achieving significantly superior open-world interaction performance compared to baseline methods.
7 Limitations
@@
Although ROCKET-1 significantly enhances interaction capabilities in Minecraft, it cannot engage with objects that are outside its field of view or have not been previously encountered. For instance, if the reasoner instructs ROCKET-1 to eliminate a sheep that it has not yet seen, the reasoner must indirectly guide ROCKET-1 ’s exploration by providing segmentations of other known objects. This limitation reduces ROCKET-1 ’s efficiency in completing simple tasks and necessitates frequent interventions from the reasoner, leading to increased computational overhead.

### v3

@@ 4.4 What Matters for Learning ROCKET-1?
We conduct ablation studies on individual tasks of Minecraft Interaction benchmark: “Hunt right sheep (
@@
Condition Fusion Methods
Fusion Positions
Hunt (
Mine (
Fusion in transformer layer
Fusion in visual backbone
Variants
#Pmt
FPS
baseline (w/o sam2)
baseline (w/o sam2)
+sam2_tiny
+sam2_small
+sam2_base_plus
+sam2_large
@@ 5 Related Works
Most current approaches (Brohan et al., 2022, 2023; Lynch et al., 2023; Cai et al., 2023a; Huang et al., 2023) use natural language to describe task details and collect large amounts of text-demonstration data pairs to train a language-conditioned policy for interaction with the environment. Although natural language can express a wide range of tasks, it struggles to represent spatial relationships effectively. Additionally, gathering text-annotated demonstration data is costly, limiting the scalability of these methods. Alternatives, such as Lifshitz et al. (2023); Majumdar et al. (2022); Sundaresan et al. (2024), use images to drive goal-conditioned policies, typically learning through hindsight relabeling in a self-supervised manner. While this reduces the need for annotated data, future images are often insufficiently expressive, making it difficult to capture detailed task execution processes. Methods like Cai et al. (2023b); Jang et al. (2022) propose using reference videos to describe tasks, offering strong expressiveness but suffering from ambiguity, which may lead to inconsistencies between policy interpretation and human understanding, raising safety concerns. Gu et al. (2023) suggests representing tasks with rough robot arm trajectories, enabling novel task completion but only in fully observable environments, limiting its applicability in open-world settings. CLIPort (Shridhar et al., 2022), which addresses pick-and-place tasks by controlling the robot’s start and end positions using heatmaps, bears some resemblance to our proposed visual-temporal context prompting method. However, CLIPort focuses solely on the pick-and-place task solutions in a fully observable environment.
@@
Minecraft offers a highly open sandbox environment with complex tasks and free exploration, ideal for testing AGI’s adaptability and long-term planning abilities. Its rich interactions and dynamic environment simulate real-world challenges, making it an excellent testbed for AGI. One line of research focuses on low-level control policies in Minecraft. Baker et al. (2022) annotated a large YouTube Minecraft video dataset with actions and trained the first foundation agent in the domain using behavior cloning, but it lacks instruction-following capabilities. Cai et al. (2023a) employs a goal-sensitive backbone and horizon prediction module to enhance multi-task execution in partially observable environments, but it only solves tasks seen during training. Fan et al. (2022) fine-tunes a vision-language alignment model MineCLIP using YouTube video data, and incorporates it into a reward shaping mechanism for training a multi-task agent, though task transfer still requires extensive environment interaction. Lifshitz et al. (2023) uses hindsight-relabeling to learn an image-goal-conditioned policy and aligns image and text spaces via MineCLIP, but this approach is limited to short-horizon tasks. Another research focus integrates vision-language models for long-horizon task planning in Minecraft (Yuan et al., 2023; Wang et al., 2024b; Qin et al., 2023; Zheng et al., 2023; Liu et al., 2024). DEPS (Wang et al., 2023b), the first to apply large language models in Minecraft, uses a four-step process to decompose tasks, achieving the diamond mining challenge with minimal training. Voyager (Wang et al., 2023a) highlights LLM-based agents’ autonomous exploration and skill-learning abilities. Jarvis-1 (Wang et al., 2023c) extends DEPS with multimodal memory, improving long-horizon task success rates by recalling past experiences. OmniJarvis (Wang et al., 2024a) learns a behavior codebook using self-supervised methods to jointly model language, images, and actions. MineDreamer (Zhou et al., 2024) fine-tunes VLMs and a diffusion model to generate goal images for task execution, though it faces challenges with image quality and consistency.
6 Conclusions and Limitations
@@
This paper presents a novel hierarchical agent architecture for open-world interaction. To address spatial communication challenges, we introduce visual-temporal context prompting to convey intent between the high-level reasoner and low-level policy. We develop ROCKET-1, an object-segmentation-conditioned policy for real-time object interaction, enhanced by SAM-2 for plug-and-play object tracking. Experiments in Minecraft show that our approach effectively leverages VLMs’ visual-language reasoning, achieving superior open-world interaction performance over baselines.
@@
Although ROCKET-1 significantly enhances interaction capabilities in Minecraft, it cannot engage with objects that are outside its field of view or have not been previously encountered. For instance, if the reasoner instructs ROCKET-1 to eliminate a sheep that it has not yet seen, the reasoner must indirectly guide ROCKET-1 ’s exploration by providing segmentations of other known objects. This limitation reduces ROCKET-1 ’s efficiency in completing simple tasks and necessitates frequent interventions from the reasoner, leading to increased computational overhead. We solve this problem in ROCKET-2 (Cai et al., 2025). This project is implemented using MineStudio (Cai et al., 2024a).
7 Acknoledgements
This work was supported by the National Science and Technology Major Project #2022ZD0114902 and the CCF-Baidu Open Fund. We sincerely appreciate their generous support, which enabled us to conduct this research.